When aqueous hydrochloric acid is poured over solid potassium sulfide, 42.9 mL of hydrogen sulfide gas is produced at a pressure of 752 torr and 25.8 °C

a. write a balanced equation for the double displacement gas-evolution reaction, including states of matter.

b. How many grams of hydrogen sulfide are produced under these parameters?
Answer #1

2HCL(aq) + K2S(s) ----> H2S(g) + 2 KCl(aq)
Lets use gas equation to get number of moles of H2S produced

PV=nRT
pressure P=752 torr=0.989 atm
Volume V= 42.9ml= 0.0429L
temperatureT= 25.8 deg = 25.8+273.15 =298.95 K
R=0.082 latm/mol/K
So number of moles
n= PV/RT = [0.989 atm*0.0429 L ]/0.082*298.95 K =0.00173 moles

molar mass of H2S=34.1 g

Mass of H2S obtained= moles* molar mass = 0.00173*34.1 g=0.0589 g

THANK YOU!!